# Causal Inference with Missing Exposures, Missing Outcomes, and Dependence

### v1

@@ Abstract
Missing data are ubiquitous in public health research. The missing-completely-at-random (MCAR) assumption is often unrealistic and can lead to meaningful bias when violated. The missing-at-random (MAR) assumption tends to be more reasonable, but guidance on conducting causal analyses under MAR is limited when there is missingness on multiple variables. We present a series of causal graphs and identification results to demonstrate the handling of missing exposures and outcomes in observational studies. For estimation and inference, we highlight the use of targeted minimum loss-based estimation (TMLE) with Super Learner to flexibly and robustly address confounding, missing data, and dependence. Our work is motivated by SEARCH-TB’s investigation of the effect of alcohol consumption on the risk of incident tuberculosis (TB) infection in rural Uganda. This study posed notable challenges due to confounding, missingness on the exposure (alcohol use), missingness on the baseline outcome (defining who was at risk of TB), missingness on the outcome at follow-up (capturing who acquired TB), and clustering within households. Application to real data from SEARCH-TB highlighted the real-world consequences of the discussed methods. Estimates from TMLE suggested that alcohol use was associated with a 49% increase in the relative risk (RR) of incident TB infection (RR=1.49, 95%CI: 1.39–1.59). These estimates were notably larger and more precise than estimates from inverse probability weighting (RR=1.13, 95%CI: 1.00–1.27) and unadjusted, complete case analyses (RR=1.18, 95%CI: 0.89–1.57). Our work demonstrates the utility of causal models for describing the missing data mechanism and TMLE for flexible inference.
@@
Keywords: Causal Inference, Missing Data, Targeted Minimum Loss-based Estimation (TMLE), Ensemble Machine Learning, Super Learner, Tuberculosis, Missing at Random
1 Introduction
Missing data affect the integrity of analyses across the spectrum of public health research, including surveillance studies to estimate disease prevalence and randomized trials to establish efficacy of new medical products [1, 2, 3, 4, 5, 6, 7, 8]. A common approach is to simply exclude observations with missing data on the relevant variables. This “complete case” analysis can provide unbiased estimates if the data are missing as a result of a completely random process (i.e., missing-completely-at-random [MCAR]) [9]. However, this assumption is often unrealistic in health settings, leading to biased estimates as well as reduced statistical power (e.g., [10]). The Missing-at-Random (MAR) assumption allows missingness to depend on observed study variables and is often more plausible [9].
Suppose, for example, we aim to estimate the prevalence of a disease in the target population. On $N$ randomly sampled participants, we measure baseline covariates (e.g., age, gender, socio-economic status), but do not fully ascertain the outcome of interest. To recover the population-level outcome prevalence under MAR, we would need that the covariates capture all the common causes of measurement and outcomes as well as a positive probability of measurement within all possible covariate values. Then using standardization (a.k.a., G-computation), inverse weighting, multiple imputation, or more robust methods, we could obtain a point estimate and 95% confidence intervals using data on all $N$ participants. Instead, if we were willing to consider the stronger MCAR assumption, corresponding to no common causes of measurement and the outcome, then we could estimate disease prevalence with an empirical proportion among those measured.
This prevalence example illustrates the consequences of causal assumptions on missing data: stronger assumptions allow for simpler estimation approaches, while weaker assumptions require more complex estimation approaches. An analogous scenario arises in analyses of observational data, which are subject to confounding. In the following, we present a comprehensive framework to address common real-world challenges: confounding, missingness on multiple variables, and dependence between participants in our study. Our work builds on a long history of methods to address missing data (e.g., [11, 9, 12, 13, 14, 15]). As detailed below, we give special consideration to missing exposures. To the best of our knowledge, methods to handle missing exposure have largely focused on case-control designs, where exposure information is retrospectively collected on a sample of participants based on their outcome status (e.g., [16, 17, 4, 18, 19, 20]). Here, we focus on standard cross-sectional or longitudinal studies with missing exposures.
@@ 1 Introduction
We also highlight the consequences of missingness on the baseline outcome when it is crucial to defining the population of interest. Suppose, for example, we are interested in studying the incidence of some disease. Our target population would be persons who are at risk of the developing the outcome and are, thereby, disease-free at baseline. In this setting, our incidence estimates would be subject to bias if there is differential measurement of outcomes at baseline. Using Counterfactual Strata Effects [21, 5, 22, 23, 24, 25, 26], we provide a framework for explicitly defining, identifying, and estimating parameters in such scenarios. Finally, we provide extensions for settings where participants are not independent and, instead, clustered within households or communities. Altogether, our presentation covers multi-source missing data, confounding, and dependence. We build-up from simple to complex examples in hope that our structured presentation is relevant a wide range of readers. We illustrate the practical relevance with the SEARCH-TB study.
2 Motivating example
@@
SEARCH was a cluster randomized trial to evaluate a community-based approach to a Universal HIV Test-and-Treat intervention, as compared to the standard-of-care, in rural Kenya and Uganda (2013-2017; NCT01864603) [27]. Following a rapid census, all communities were offered multi-disease testing through community health campaigns with home-based follow-up for non-participants [28]. Through this mechanism, we measured demographic data (e.g., age, sex, education, and mobility), self-reported alcohol use (our primary exposure of interest), and tested for HIV infection. Due to high costs and complex logistics, evaluation of incident tuberculosis (TB) infection, a proxy for population-level transmission, was limited to SEARCH-TB, a sub-study in 9 eastern Ugandan communities [29, 30]. This sub-study was intentionally enriched for persons with HIV. Specifically, in each community, we sampled 100 households with at least one adult (15+ years) with HIV and 100 households without an adult with HIV. At baseline of the sub-study, tuberculin skin tests were administered to residents of the sampled households. One year later, follow-up tests were administered to participants who tested negative at baseline. Here, we provide an in-depth exploration of the methods to evaluate the effect of alcohol use on incident TB infection in SEARCH-TB [31]: (1) missingness on the exposure of interest, (2) missingness on the baseline outcome, crucial to defining the target population of interest, (3) missingness in the final outcome, and (4) confounding. Additionally, given our focus on infectious disease outcomes, the independent and identically distributed (i.i.d.) assumption was likely violated, and we accounted dependence among study participants.
3 Related Causal Problems: Building Complexity
@@
Many studies feature only a subset of the challenges described above. We, thus, provide causal models and identification results for a series of hypothetical studies with increasing complexity in the hope of providing a useful reference for a broad range of real-world studies. For simplicity, we focus on defining and identifying causal parameters under a single level of the exposure, but our results naturally generalize to causal effects defined in terms of contrasts of counterfactual outcome distributions under two levels of the exposure (i.e., the average treatment effect or the causal risk ratio).
3.1 Classic point-treatment problem
First, we consider the classic “point-treatment” problem, where we have measured confounding by baseline covariates $L$, a binary exposure $A$ occurring at single time-point, and an outcome $Y$ occurring at the study’s close. This could represent a study of the effect of alcohol use ($A$) on incident TB infection ($Y$) among a representative cohort of persons without TB at baseline. The directed acylic graph (DAG) and non-parametric structural equation model (NPSEM) for such a study are given in Figure 1 and reflect the simplifying assumption that participants are independent (e.g., cannot transmit TB to one another due to geographical distance).
Figure 1: Directed acyclic graph (DAG) for a classic point-treatment problem with complete measurement of the baseline covariates $L$, the exposure $A$, and the outcome $Y$. The corresponding non-parametric structural equation model (NPSEM) is given by $L=f_{L}(U_{L})$; $A=f_{A}(L,U_{A})$; $Y=f_{Y}(L,A,U_{Y})$ where $(U_{L},U_{A},U_{Y})$ represent unmeasured factors determining the values of the covariates, exposure, and outcome, respectively. On the DAG, $U^{*}$ represents unmeasured common causes of at least two variables in $(L,A,Y)$.
To identify our causal target parameter and express it as function of the distribution of the observed data $O=(L,A,Y)$, we would need there to be no unmeasured confounding, which corresponds to the assumption that the baseline covariates $L$ capture all the joint causes of the exposure $A$ and outcome $Y$ and can be represented as $Y^{a}\perp A\mid L$. Additionally, we need there to be a non-zero probability of having the exposure in all possible values of $L$: $\mathbb{P}(A=a\mid L)>0$ a.e.. Under these two assumptions, our causal target is equal to the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$ [12]. Even if these assumptions are not reasonable (e.g., there are unmeasured confounders in Figure 1), we still have well-defined statistical estimand, on which we can focus our estimation efforts.
3.2 Missing exposures
@@
Most real-world studies, however, depart from this idealized point-treatment problem. The first departure we consider is missingness on the exposure of interest. Continuing our running example, suppose that among our representative cohort of persons without TB at baseline, some participants did not answer questions about their alcohol use. Let $\Delta_{A}$ be an indicator that a participant has their exposure measured. If $\Delta_{A}=1$ for a participant, we observe their exposure $A$ as usual. However, if $\Delta_{A}=0$ for a participant, their exposure $A$ is not observed (i.e., equal to NA). The DAG and NPSEM for such a study are given in Figure 2. To define causal effects when the exposure is subject to missingness, we now consider counterfactuals indexed by both the exposure and its measurement indicator. Specifically, let $Y^{*}=Y^{\Delta_{A}=1,A=a}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, their exposure were measured and was at level $A=a$.
Figure 2: DAG for a point-treatment problem with missingness on the exposure: $L$=baseline covariates, $\Delta_{A}$=indicator of exposure measurment, $A$=exposure, $Y$=outcome, and $U^{*}$=unmeasured common causes of at least two variables in $(L,\Delta_{A},A,Y)$. The corresponding NPSEM is $L=f_{L}(U_{L})$; $\Delta_{A}=f_{\Delta_{A}}(L,U_{\Delta_{A}})$, $A=f_{A}(L,\Delta_{A},U_{A})$, $Y=f_{Y}(L,\Delta_{A},A,U_{Y})$.
Then our causal target parameter $\mathbb{E}(Y^{*})$ is the counterfactual mean outcome if all participants had their exposure measured $\Delta_{A}=1$ and it were at level $A=a$. To identify this causal parameter and express it as a function of the distribution of the observed data $O=(L,\Delta_{A},A,Y)$, we would need $L$ to be sufficient to control for confounding and differential measurement. In more detail, the following assumptions are required:
  • Within values of the covariates $L$, outcomes among those with measured exposures are representative of outcomes among those without measured exposures: $Y^{*}\perp\Delta_{A}\mid L$.
  • There is a non-zero probability of exposure measurement within all possible values of $L$: $\mathbb{P}(\Delta_{A}=1\mid L)>0$ a.e..
  • Among those with measured exposures ($\Delta_{A}=1$),
    • $L$ captures all the common causes of the exposure and outcome: $Y^{*}\perp A\mid\Delta_{A}=1,L$.
    • there is a non-zero probability of being exposed to level $A=a$ within all possible values of $L$: $\mathbb{P}(A=a\mid\Delta_{A}=1,L)>0$ a.e..
If these assumptions hold, we can rewrite $\mathbb{E}(Y^{*})$ as the statistical estimand $\mathbb{E}[\mathbb{E}(Y\mid A=a,\Delta_{A}=1,L)]$ with proof in eAppendix A. As before, even if these assumptions do not hold (e.g., there are unmeasured common causes of measurement and the outcome in Figure 2), we still have a well-defined statistical estimand for estimation and inference.
3.3 Missing Exposures and Outcomes
@@
We now add another common complication: missing outcomes. Continuing our running example, suppose that among our cohort of persons at risk of TB, some participants did not answer questions about their alcohol use and, despite best efforts, some participants could not be found at the end of the study for outcome ascertainment. To reflect this data generating process, we introduce new notation to reflect the longitudinal data setting. Let $L_{0}$ be baseline covariates, $L_{1}$ be additional covariates collected after the exposure but before outcome ascertainment, and $\Delta_{Y}$ be an indicator of outcome measurement. Specifically, if $\Delta_{Y}=1$ for a participant, we observe their outcome $Y$ as before. However, if $\Delta_{Y}=0$ for a participant, their outcome $Y$ is not observed (i.e., equal to NA). The simplified DAG and full NPSEM for such a study are given in Figure 3.
Figure 3: DAG with missingness on the exposure and outcome: $L_{0}$=baseline covariates, $\Delta_{A}$=indicator of measured exposure, $A$=exposure, $L_{1}$=time-varying covariates, $\Delta_{Y}$=indicator of measured outcome, and $Y$=outcome. For simplicity, we have omitted the $U^{*}$ node, representing unmeasured common causes of at least two variables in $(L_{0},\Delta_{A},A,L_{1},\Delta_{Y},Y)$. The NPSEM is given by $L_{0}=f_{L_{0}}(U_{L_{0}})$; $\Delta_{A}=f_{\Delta_{A}}(L_{0},U_{\Delta_{A}})$; $A=f_{A}(L_{0},\Delta_{A},U_{A})$; $L_{1}=f_{L_{1}}(L_{0},\Delta_{A},A,U_{L_{1}})$; $\Delta_{Y}=f_{\Delta_{Y}}(L_{0},\Delta_{A},A,L_{1},U_{\Delta_{Y}})$; $Y=f_{Y}(L_{0},\Delta_{A},A,L_{1},\Delta_{Y},U_{Y})$.
@@
To define the causal effect when the exposure and outcome are subject to missingness, we now consider counterfactuals indexed by the exposure and two measurement indicators. Specifically, let $Y^{*}=Y^{\Delta_{A}=1,A=a,\Delta_{Y}=1}$ denote the counterfactual outcome under hypothetical interventions to ensure exposure measurement, “set” the exposure level to $A=a$, and ensure outcome measurement. To identify the counterfactual mean outcome $\mathbb{E}(Y^{*})$ and express it as a function of the distribution of the observed data $O=(L_{0},\Delta_{A},A,L_{1},\Delta_{Y},Y)$, we now need to account for the post-baseline covariates $L_{1}$, which act as time-dependent confounders. Specifically, $L_{1}$ are mediators of the exposure-outcome relationship, while “confounding” the measurement-outcome relationship. Therefore, we rely on sequential randomization and find a set of covariates that satisfies the backdoor criteria for each “intervention” node given the observed past [12]. As before, we need that the baseline covariates $L_{0}$ are sufficient to control for missing exposures and for confounding. In other words, we need the analogous identification assumptions given in the prior subsection. Additionally, we need that among participants with measured exposures at the level of interest (i.e., $\Delta_{A}=1$ and $A=a$),
  • the baseline and time-varying covariates ($L_{0},L_{1})$ capture all the common causes of outcomes and their measurement: $Y^{*}\perp\Delta_{Y}\mid L_{1},A=a,\Delta_{A}=1,L_{0}$.
  • there is a positive probability of outcome measurement within all possible values of the baseline and time-varying covariates: $\mathbb{P}(\Delta_{Y}=1\mid L_{1},A=a,\Delta_{A}=1,L_{0})>0$ a.e..
If these assumptions hold, we can rewrite $\mathbb{E}(Y^{*})$ in terms of the longitudinal G-computation formula: $\mathbb{E}\{\mathbb{E}[\mathbb{E}(Y\mid\Delta_{Y}=1,L_{1},A=a,\Delta_{A}=1,L_{0})\mid A=a,\Delta_{A}=1,L_{0})]\}$, shown in terms of iterated expectations and with proof in eAppendix B [12, 32, 33]. As before, even if these identification assumptions do not hold, we still have a well-defined statistical estimand to focus our estimation efforts.
3.4 Missing Exposures and Outcomes at Baseline and Follow-up
@@
We now consider missingness on the outcome at baseline. Continuing our running example, we are interested in the effect of alcohol use on incident TB infection, but did not reach 100% of study participants for baseline TB testing. In other words, our cohort of participants without TB at baseline is subject to selection bias. To reflect this data generating process, we update our notation to have multiple outcome measures. Let $\Delta_{Y_{0}}$ be an indicator of outcome measurement at baseline, $Y_{0}$ be an indicator having the outcome at baseline, $\Delta_{Y_{1}}$ be an indicator of outcome measurement at follow-up, and $Y_{1}$ be an indicator having the outcome at follow-up. As previously introduced, the value of the baseline outcome defines the target population; specifically, we are interested in occurrence of the outcome ($Y_{1}=1$) among those who are at risk at baseline ($Y_{0}=0$). The corresponding causal models can be found in Figure 4.
Figure 4: DAG with missingness on the exposure, the baseline outcome, and the follow-up outcome: $L_{0}$=baseline covariates, $\Delta_{A}$=indicator of measured exposure, A=exposure, $\Delta_{Y_{0}}$=indicator of measured baseline outcome, $Y_{0}$=baseline outcome, $L_{1}$=time-dependent covariates, $\Delta_{Y_{1}}$=indicator of measured outcome at follow-up, and $Y_{1}$= outcome at follow-up. For simplicity, we have omitted the $U^{*}$ node, representing unmeasured common causes of at least 2 variables in $(L_{0},\Delta_{A},A,\Delta_{Y_{0}},Y_{0},L_{1},\Delta_{Y_{1}},Y_{1})$. The NPSEM is given by $L_{0}=f_{L_{0}}(U_{L_{0}})$; $\Delta_{A}=f_{\Delta_{A}}(L_{0},U_{\Delta_{A}})$; $A=f_{A}(L_{0},\Delta_{A},U_{A})$; $\Delta_{Y_{0}}=f_{\Delta_{Y_{0}}}(L_{0},\Delta_{A},A,U_{\Delta_{Y_{0}}})$; $Y_{0}=f_{Y_{0}}(L_{0},\Delta_{A},A,\Delta_{Y_{0}},U_{Y_{0}})$; $L_{1}=f_{L_{1}}(L_{0},\Delta_{A},A,\Delta_{Y_{0}},Y_{0},U_{L_{1}})$; $\Delta_{Y_{1}}=f_{\Delta_{Y_{1}}}(L_{0},\Delta_{A},A,\Delta_{Y_{0}},L_{1},Y_{0},U_{\Delta_{Y_{1}}})$; $Y_{1}=f_{Y_{1}}(L_{0},\Delta_{A},A,\Delta_{Y_{0}},L_{1},Y_{0},\Delta_{Y_{1}},U_{Y_{1}})$.

### v2

# Causal Inference with Missing Exposures and Missing Outcomes

@@ Abstract
Missing data are ubiquitous in public health research. When estimating causal effects, there are well-established methods to address bias to due censored outcomes. Commonly, causal estimands are defined under hypothetical interventions to “set” the exposure and to prevent censoring. Identification is evaluated with the sequential backdoor criterion and considerations of data support. Then inverse weighting, standardization, and doubly-robust approaches are applied for statistical estimation and inference. We demonstrate how this framework can be extended to settings with missingness on the exposure of interest as well as the variable defining the population of interest (e.g., persons at risk of the outcome). Our work is motivated by SEARCH-TB’s investigation of the effect of alcohol consumption on the risk of incident tuberculosis (TB) infection in rural Uganda. This study posed several real-world challenges: confounding, missingness on the exposure (alcohol use), missingness on the baseline outcome (defining who was at risk of TB), and missingness on the outcome at follow-up (capturing who acquired TB). We present a series of causal models and identification results to demonstrate the handling of missing exposures and outcomes in prospective studies. We highlight the use of TMLE with Super Learner and the real-world consequences of our approach.
@@
Keywords: Causal Inference, Missing Data, Real-world Evidence, Super Learner, Targeted Minimum Loss-based Estimation, TMLE, Tuberculosis
Introduction
Missing data affect the integrity of analyses across the spectrum of public health research, including surveillance studies to estimate disease prevalence and randomized trials to establish efficacy of new medical products [1, 2, 3, 4, 5, 6, 7, 8]. There is rich history of methods research to address the potential for bias when participants with measured outcomes differ meaningfully from those with missing or censored outcomes (e.g., [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]). The Causal Roadmap provides one such approach [21, 22, 23]. First, we specify our causal question: how would expected outcomes differ if all did versus did not receive the exposure of interest and censoring were prevented. Then, we specify a causal model, such as directed acyclic graph (DAG) or non-parametric structural equation model (NPSEM), to represent the data generating process for observed data: baseline and time-varying confounders, exposures, and outcomes [10]. Third, we intervene on the causal model to generate counterfactual outcomes under hypothetical interventions to set the exposures and ensure outcome measurement. Fourth, we evaluate whether the corresponding causal estimand can be identified (i.e., if sequential exchangeability and positivity hold) [9, 10, 11, 12]. Fifth, we specify the statistical estimand, which is often a complex function of the observed data distribution (i.e., not equal to single regression coefficient). Sixth, we conduct estimation and inference with inverse weighting, standardization, or doubly-robust methods, such as targeted minimum loss-based estimation (TMLE) [11]. Finally, we conduct sensitivity analyses and appropriately interpret the results.
@@ Introduction
Here, we provide a tutorial on how this framework can be extended to address missingness on the exposure. We also highlight the consequences of missingness on the baseline outcome when it is crucial to defining the population of interest. Suppose, for example, we are interested in studying the incidence of some disease. Our population of interest would be persons who are at risk of the developing the outcome and are, thereby, disease-free at baseline. In this setting, our incidence estimates would be subject to bias if there is differential missingness of outcomes at baseline. Using Counterfactual Strata Effects [24, 5, 25, 26, 27, 28, 29], we provide a framework for explicitly defining, identifying, and estimating parameters in such scenarios.
We note there is a growing interest in the use of missingness graphs to represent studies with missingness on multiple variables and to assess whether causal effects can be identified (termed “recovered”) [30, 31]. In particular, Moreno-Betancur and colleagues introduced complete-data DAGs (c-DAGs), missingness DAGs (m-DAGs), and their link [4]. They also provided a series of “canonical” m-DAGs for causal effects of point-treatment exposures [4, 32]. Holovchak et al. extended this work for the effects of longitudinal exposures [33], while questioning the practical utility of the m-DAG approach. Specifically, they noted, “no general algorithms are available to decide on recoverability, and decisions have to be made on a case-by-case basis” [33]. As an alternative, we demonstrate how an established framework for causal inference can used to define, identify, and estimate causal estimands, such as the average treatment effect, under missingness on exposures and outcomes. In the following, we build-up from simple to complex examples in hope that our structured presentation is relevant a wide range of readers who need to address confounding and multi-source missingness. We illustrate practical relevance with the SEARCH-TB study.
Motivating example
@@
SEARCH was a cluster randomized trial to evaluate a community-based approach to a Universal HIV Test-and-Treat intervention, as compared to the standard-of-care, in rural Kenya and Uganda (2013-2017; NCT01864603) [34]. Following a rapid census, all communities were offered multi-disease testing through community health campaigns with home-based follow-up for non-participants [35]. Through this mechanism, we measured demographic data (e.g., age, sex, education, and mobility), self-reported alcohol use (our primary exposure of interest), and tested for HIV infection. Due to high costs and complex logistics, evaluation of incident tuberculosis (TB) infection was limited to SEARCH-TB, a sub-study in 9 eastern Ugandan communities [36, 37]. This sub-study was intentionally enriched for persons with HIV. Specifically, in each community, we sampled 100 households with at least one adult (15+ years) with HIV and 100 households without an adult with HIV. At baseline of the sub-study, tuberculin skin tests were administered to residents of the sampled households. One year later, follow-up tests were administered to participants who tested negative at baseline. Here, we demonstrate the methods used to evaluate the effect of alcohol use on incident TB infection [38]: confounding, missingness on the exposure of interest, missingness on the baseline outcome (defining who was at risk of TB), and missingness on the final outcome (defining who acquired TB).
Related Causal Problems: Building Complexity
@@
Many studies feature only a subset of the challenges described above. We, thus, provide causal models and identification results for a series of hypothetical studies with increasing complexity in the hope of providing a useful reference for a broad range of real-world studies. For simplicity, we focus on defining and identifying causal parameters under a single level of the exposure, but our results naturally generalize to causal effects defined in terms of contrasts of counterfactual outcome distributions under two levels of the exposure (i.e., the average treatment effect or causal risk ratio).
Classic point-treatment problem
In Appendix S1, we review the classic “point-treatment” problem, where we have measured confounding by baseline covariates $L$, a binary exposure $A$ occurring at single time-point, and an outcome $Y$ occurring at the study’s close. This could represent a study of the effect of alcohol use ($A$) on incident TB infection ($Y$) among a representative cohort of persons without TB at baseline. Let $Y^{a}$ be the counterfactual outcome if, possibly contrary-to-fact, the participant had exposure-level $A=a$. The counterfactual mean outcome $\mathbb{E}(Y^{a})$ is identified if there are no unmeasured confounders and sufficient data support: $Y^{a}\perp A\mid L$ and $\mathbb{P}(A=a\mid L)>0$, respectively. The statistical estimand is given by the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$[9]. Even if these assumptions are not reasonable, the G-computation formula is still a well-defined statistical estimand, on which we can focus our estimation efforts (Appendix S1).
Missing Outcomes
Most real-world studies, however, depart from this idealized point-treatment problem. Here, we review causal models and identification when there is missingness on the outcome. Continuing our running example, suppose that among our representative cohort of persons without TB at baseline, some participants did not test for TB at the end of follow-up. Let $\Delta_{Y}$ be an indicator of outcome measurement. If $\Delta_{Y}=1$ for a participant, we observe their outcome $Y$ as usual. However, if $\Delta_{Y}=0$, their outcome $Y$ is not observed. In Figure 1, we show two possible ways of using causal models to represent such a study. In Panel A, we introduce $Y^{\circ}$ to represent the underlying value of the outcome, and define the observed outcome as the product of its measurement indicator and underlying value: $Y=\Delta_{Y}\times Y^{\circ}$. Thus, $Y$ equals its underlying value $Y^{\circ}$ when it’s measured (i.e., when $\Delta_{Y}=1$) and is zero otherwise (i.e., when $\Delta_{Y}=0$). In Panel B, we omit the underlying outcome and directly represent the causal model in terms of the observed data: $O=(L,A,\Delta_{Y},Y)$. In the latter presentation, the measurement indicator $\Delta_{Y}$ again influences the value of the observed outcome $Y$, because when an outcome is not measured, it is set deterministically to NA. Both approaches lead us to same identification assumptions and statistical estimands. Therefore, to minimize notation and mirror the long-standing literature on censoring (e.g., [16, 17, 11, 18, 19, 20, 12]), we use the latter representation in the remainder of the article.
To define causal effects when the outcome is subject to missingness, we now consider counterfactuals indexed by both the exposure and the outcome measurement indicator. Specifically, let $Y^{*}=Y^{A=a,\Delta_{Y}=1}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, their exposure were at level $A=a$ and missingness on the outcome prevented. Then our causal target parameter $\mathbb{E}(Y^{*})$ is the counterfactual mean outcome if all participants had exposure $A=a$ and their outcomes measured. To identify this causal parameter and express it as a function of the observed data distribution, we would need the baseline covariates $L$ to be sufficient to control for confounding and differential outcome missingness. Specifically, we need $L$ to capture all the common causes of the exposure and outcome, and, among those with the exposure of interest, all the common causes of the outcome and its measurement. These assumptions can be represented as $Y^{*}\perp A\mid L$ and $Y^{*}\perp\Delta_{Y}\mid A=a,L$, respectively. These conditions are often referred to as the “sequential randomization assumption” or “sequential exchangeability” and can be evaluated graphically through the sequential backdoor criterion [9, 10, 11, 12]. We also need a positive probability of being exposed to level $A=a$ within all possible values of the confounders and, among those with the exposure of interest, a positive probability of outcome measurement within all possible values of the confounders: $\mathbb{P}(A=a\mid L)>0$ a.e. and $\mathbb{P}(\Delta_{Y}\mid A=a,L)>0$ a.e., respectively. If these assumptions hold, we have equivalence between our causal estimand $\mathbb{E}(Y^{*})$ and the statistical estimand given by the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid\Delta_{Y}=1,A=a,L)]$ [9].
Figure 1: Causal models with missingness on the outcome (Panels A-B) or missingness on the exposure (Panel C): LL=baseline covariates, $\Delta_{A}$=indicator of exposure measurement, AA=observed exposure, $Y^{\circ}$=underlying outcome, $\Delta_{Y}$=indicator of outcome measurement, YY= observed outcome. For ease of presentation, the models are shown without dependence between the unmeasured variables, which are omitted on the graph.
Missing Exposures
@@
We now demonstrate how the above approach for missing outcomes can be applied to missing exposures. Continuing our running example, suppose that among our representative cohort of persons without TB at baseline, some participants did not answer questions about their alcohol use. Let $\Delta_{A}$ be an indicator that a participant has their exposure measured. If $\Delta_{A}=1$ for a participant, we observe their exposure $A$ as usual. However, if $\Delta_{A}=0$ for a participant, their exposure $A$ is set deterministically to NA. The DAG and NPSEM for such a study are given in Figure 1C. To define causal effects when the exposure is subject to missingness, we consider counterfactuals indexed by both the exposure and its measurement indicator. Now, let $Y^{*}=Y^{\Delta_{A}=1,A=a}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, their exposure were measured and at level $A=a$.
Then to identify the expected counterfactual outcome $\mathbb{E}(Y^{*})$ and express it as a function of the distribution of the observed data $O=(L,\Delta_{A},A,Y)$, we need analogous conditions as the prior subsection on missing outcomes. Specifically, we need $L$ to be sufficient to control for differential exposure missingness and for confounding (among those with measured exposures): $Y^{*}\perp\Delta_{A}\mid L$ and $Y^{*}\perp A\mid\Delta_{A}=1,L$, respectively. These assumptions can be graphically evaluated with the sequential backdoor criterion. We also need the two analogous assumptions on data support: $\mathbb{P}(\Delta_{A}=1\mid L)>0$ a.e. and $\mathbb{P}(A=a\mid\Delta_{A}=1,L)>0$ a.e. If these four assumptions hold, we can rewrite $\mathbb{E}(Y^{*})$ as the statistical estimand $\mathbb{E}[\mathbb{E}(Y\mid A=a,\Delta_{A}=1,L)]$ with proof in Appendix S2.1.
Missing Exposures and Outcomes
@@
We now combine our two challenges: missing exposures and missing outcomes. Continuing our running example, suppose that among our cohort of persons at risk of TB, some participants did not answer questions about their alcohol use and, despite best efforts, some participants could not be found at the end of the study for outcome ascertainment. To reflect this data generating process, we introduce new notation to reflect the longitudinal data setting. Let $L_{0}$ be baseline covariates and $L_{1}$ be additional covariates collected after the exposure but before outcome ascertainment. Figure 2 provides the causal models for such a study.
Figure 2: Causal graph and corresponding non-parametric structural equation model with missingness on the exposure and the outcome: $L_{0}$=baseline covariates, $\Delta_{A}$=indicator of exposure measurement, AA=observed exposure, $L_{1}$=time-varying covariates, $\Delta_{Y}$=indicator of outcome measurement, and YY=observed outcome. For ease of presentation, the models are shown without dependence between the unmeasured variables, which are omitted on the graph.
@@
To define the causal effect when the exposure and outcome are subject to missingness, we now consider counterfactuals indexed by the exposure and two measurement indicators. Specifically, let $Y^{*}=Y^{\Delta_{A}=1,A=a,\Delta_{Y}=1}$ denote the counterfactual outcome under hypothetical interventions to ensure exposure measurement, “set” the exposure level to $A=a$, and ensure outcome measurement. To identify the counterfactual mean outcome $\mathbb{E}(Y^{*})$ and express it as a function of the distribution of the observed data $O=(L_{0},\Delta_{A},A,L_{1},\Delta_{Y},Y)$, we now need to account for the post-baseline covariates $L_{1}$, which act as time-dependent confounders. Specifically, $L_{1}$ are mediators of the exposure-outcome relationship, while “confounding” the measurement-outcome relationship. Therefore, we rely on sequential randomization/exchangeability and find a set of covariates that satisfies the backdoor criterion for each “intervention” node given the observed past [9]. As before, we need that the baseline covariates $L_{0}$ are sufficient to control for missing exposures and for confounding. In other words, we need the analogous identification assumptions given in the prior subsection. Additionally, we need that among participants with measured exposures at the level of interest (i.e., $\Delta_{A}=1$ and $A=a$), the baseline and time-varying covariates ($L_{0},L_{1})$ capture all the common causes of outcomes and their measurement as well as a positive probability of outcome measurement within all possible values of the baseline and time-varying covariates: $Y^{*}\perp\Delta_{Y}\mid L_{1},A=a,\Delta_{A}=1,L_{0}$ and $\mathbb{P}(\Delta_{Y}=1\mid L_{1},A=a,\Delta_{A}=1,L_{0})>0$ a.e., respectively. If these assumptions hold, we can rewrite $\mathbb{E}(Y^{*})$ in terms of the longitudinal G-computation formula: $\mathbb{E}\{\mathbb{E}[\mathbb{E}(Y\mid\Delta_{Y}=1,L_{1},A=a,\Delta_{A}=1,L_{0})\mid A=a,\Delta_{A}=1,L_{0})]\}$, shown in terms of iterated expectations and with proof in Appendix S2.2 [9, 39, 40].
Missing Exposures and Missing Outcomes at the Start and End of Follow-up
@@
We now consider missingness on the outcome at the start of follow-up. Following our motivating example, we are interested in the effect of alcohol use on incident TB infection, but did not reach 100% of study participants for baseline TB testing. In other words, our longitudinal cohort of participants without TB is subject to selection bias. To reflect this data generating process, we update our notation to have multiple outcome measures. Let $\Delta_{Y_{0}}$ be an indicator of outcome measurement at start of follow-up (hereafter “baseline”) and $\Delta_{Y_{1}}$ be an indicator of outcome measurement at the end of follow-up (hereafter “endline”). Let $Y_{0}$ and $Y_{1}$ denote the corresponding outcomes; they are set to NA when not observed. The corresponding causal models can be found in Figure 3. The time-ordering reflects the study protocol and procedures in SEARCH-TB; specifically, alcohol use $A$ was measured before baseline TB status $Y_{0}$. As a result, the exposure $A$ can impact who has prevalent TB at baseline ($Y_{0}=1$) and, thereby, who is at risk of TB at baseline ($Y_{0}=0$).
Figure 3: Causal graph and corresponding non-parametric structural equation model with missingness on the exposure and the outcome at baseline and endline: $L_{0}$=baseline covariates, $\Delta_{A}$=indicator of exposure measurement, A=exposure, $\Delta_{Y_{0}}$=indicator of outcome measurement at baseline, $Y_{0}$=baseline outcome, $L_{1}$=time-dependent covariates, $\Delta_{Y_{1}}$=indicator of outcome measurement at endline, and $Y_{1}$= outcome at endline. For ease of presentation, the models are shown without dependence between the unmeasured variables, which are omitted on the graph.